OpenStudy (anonymous):

OK here is a stumper... Solve for b: 5b^2-5b+1=0

OpenStudy (anonymous):

Solve for b: \[5b ^{2}-5b+1\]

OpenStudy (anonymous):

\[5+\sqrt{5}/10, 5-\sqrt{5}/10\]

OpenStudy (anonymous):

1+sqrt(5)/2 0r 1-sqrt(5)/2

OpenStudy (anonymous):

by using quadratic formula \[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] where a=5,b=-5 and c=1 plug the values into the formula \[\frac{-(-5)\pm \sqrt{(-5)^2-4(5)(1)}}{2(5)}\] \[\frac{5\pm \sqrt{25-20}}{10}\] \[\frac{5\pm \sqrt{5}}{10}\] \[\frac{5+\sqrt{5}}{10},\frac{5-\sqrt{5}}{10}\]

OpenStudy (anonymous):

But the simplified answer is what I gave

OpenStudy (anonymous):

yep that was it!!!! thanks